# Josephson effect and self-trapping in helicoidal spin-orbit coupled Bose-Einstein condensates with optical lattices

### v3

@@ I Introduction
Josephson physics describes the tunnelling phenomena that arise due to coupling between two macroscopically phase coherent quantum states such as superconductors separated by thin insulating layer and, weakly coupled Bose-Einstein condensates (BEC) separated by tunnelling layer. The interplay between tunnelling and interaction results supercurrent due to the oscillation of population difference between these two states [1, 2]. This is the so-called Josephson oscillation. A transition from Josephson oscillation to self-trapping occurs if interaction dominates over the tunnelling and causes the population imbalance to oscillate about a non-zero value. This tunnelling oscillation can be explored in building quantum mechanical circuit to design qubit [3], SQUIDS [4] and to engineer quantum states [5].
Two Bose-Einstein condensates in a trap consisting of atoms differing in hyperfine states provide a platform to realize bosonic Josephson junction of a weakly interacting system. A spin-orbit coupled (SOC) atomic BEC created by artificial gauge potentials offers a new possibility to study Josephson oscillation in pseudo-spin 1/2 systems. This type of oscillation in a weakly interacting BEC is a nonlinear generalization of the typical a.c. and d.c. Josephson oscillations. Thus it has received special attention and resulted several interesting experimental [6, 7, 8] and theoretical [9, 10, 11, 12, 13, 14, 15, 16, 17] investigations. Particularly, theoretical investigations in this direction include the studies on Josephson dynamics in two species BECs in double-well potential [10], internal Josephson oscillation between coupled solitons with time varying Raman frequency [17], effects of SOC on Josephson tunnelling between two annular condensates [18], Josephson oscillation of 2D chiral solitons [19] and Josephson-type oscillation in presence of optical lattices [20].
@@
Various techniques for generating artificial gauge potentials allow to tune strength and shape of spin-orbit coupling in BECs [26, 27, 28]. For the choice of gauge potential, $A(x)=\sigma n(x)$, where $n(x)=(\cos(2\beta x),\sin(2\beta x),0)$ and $\sigma=(\sigma_{1},\sigma_{2},\sigma_{3})$, one can induce helicoidal shaped non-uniform spin-orbit coupling in BECs. This non-uniform SOC is highly tunable and can increase the effects of intrinsic nonlinearity [26, 27, 28, 29, 30, 31]. A Bose-Einstein condensate with helicoidal SOC features many properties of integrable system such as stable propagation of bright solitons [32], Peregrine solitons [33], bright-dark Peregrine solitons [34], stripe solitons [35], multi-hump solitons [36] and kink solitons [37]. One of the important aspects of the soliton in helicoidal SOC-BECs is that the soliton with any peak amplitude is non-radiative and thus can propagate a large distance [32]. In a recent study it is shown that the helicoidal gauge potential can be used to manipulate modulational instability in BECs [38].
@@
Our objective in this paper is to study Josephson-type (JT) oscillation and self-trapping of atomic population between two solitons in helicoidal spin-orbit coupled quasi-one-dimensional (Q1D) BECs with optical lattices. Particularly, we see that the atomic imbalance between the two solitons oscillates periodically and the frequency of oscillation depends sensitively on the strength of helicoidal gauge potential. The frequency decreases with the increase of helicoidal gauge potential. In presence of optical lattices, the frequency of oscillation further increases [20]. We see that there exists a critical point for the observation of self-trapping. Interestingly, the critical point gets shifted in the parameter space with the change of helicoidal gauge potential and Zeeman-splitting. We note that there is no Rabi coupling in helicoidal SOC-BEC. Therefore, the oscillation of population imbalance occurs due to the helicoidal SOC. In the case of a traditional SOC-BEC, the states are interacting through Rabi coupling [39, 40, 41].
@@
In section II, we present a theoretical model to describe helicoidal SOC-BEC in optical lattices. Based on a variational approach, we derive equations for the center of mass, population imbalance and phase difference of the coupled condensates. In section III, we study the dynamics of population imbalance for a particular location of center of mass in the linar optical lattice. Specifically, we study the effects of nonlinear optical lattices on JT oscillations for different strengths of the helicoidal gauge potentials. In section IV, we find a condition for the observation of quantum mechanical self-trapping and its dependence on the helicoidal gauge potential and nonlinear optical lattice. Finally, we devote section V to summarize the results of our investigations.
@@ II General Formalism
$i\frac{\partial\psi_{1}}{\partial t}$ $=$ $-\frac{1}{2}\frac{\partial^{2}\psi_{1}}{\partial x^{2}}+i\beta\frac{\partial\psi_{1}}{\partial x}-i\alpha_{s}\frac{\partial\psi_{2}}{\partial x}+V_{ext}(x)\psi_{1}$ (1)
$+$ $\frac{\Delta}{2}\psi_{1}-(g_{1}|\psi_{1}|^{2}+g_{2}|\psi_{2}|^{2})\psi_{1},$
$i\frac{\partial\psi_{2}}{\partial t}$ $=$ $-\frac{1}{2}\frac{\partial^{2}\psi_{2}}{\partial x^{2}}-i\beta\frac{\partial\psi_{2}}{\partial x}-i\alpha_{s}\frac{\partial\psi_{1}}{\partial x}+V_{ext}(x)\psi_{2}$ (2)
-- $\frac{\Delta}{2}\psi_{2}-(g_{1}|\psi_{2}|^{2}+g_{2}|\psi_{1}|^{2})\psi_{2}$
@@
$V_{ext}=\frac{1}{2}{({\omega_{x}}/{\omega_{\perp}})^{2}}x^{2}+V_{0}\cos(2k_{l}x).$ (3)
Here, $\psi_{j}(x,t)$ $(j=1,2)$ gives the order parameter. The strengths of inter- and intra-atomic interactions are represented by $g_{1}$ and $g_{2}$. In Eqs. (1) and (2), the parameters $\alpha_{s}$, $\beta$ and $\Delta$ stand for the strengths of spin-orbit coupling, helicoidal gauge potential (HGP) and Zeeman splitting respectively. The external potential in Eq. (3) contains two terms, namely, harmonic trapping potential (first term) and linear optical lattice (second term) with wavenumber $k_{l}$ and strength $V_{0}$. We make Eqs. (1) and (2), dimensionless by re-scaling the energy, length and time of the system in the units of $\hbar\omega_{\perp}$, $a_{\perp}=\sqrt{{\hbar}/{m\omega_{\perp}}}$ and $({\omega_{\perp}})^{-1}$ respectively. Here $\omega_{x}$ and $\omega_{\perp}$ stand for the angular frequencies of the harmonic trap in the axial and radial directions.
With a view to consider a helicoidal spin-orbit coupled BEC trapped in linear optical lattices (LOL), we take $\omega_{\perp}\gg\omega_{x}$ such that the effects of harmonic trap is negligible. In addition, we introduce periodicity in the system by modulating the nonlinear mean-field interaction periodically, called nonlinear optical lattice (NOL). The NOLs are given by
@@
where $k_{n}$ is the wavenumber of the NOLs with their strengths $\gamma_{1}$ and $\beta_{1}$ respectively. The NOLs are found very effective in controlling the dynamics and stability of matter-wave solitons in BECs [42, 43, 44, 45, 46, 47]. In order to proceed with the variational approach, we recast the Eqs. (1) and (2) in the following Lagrangian density.
$\mathcal{L}$ $=$ $\bigg{[}\frac{i}{2}\bigg{(}\psi_{1}^{*}\frac{d\psi_{1}}{dt}+\psi_{2}^{*}\frac{d\psi_{2}}{dt}\bigg{)}$ (5)
-- $\frac{i\beta}{2}\bigg{(}\psi_{1}^{*}\frac{d\psi_{1}}{dx}-\psi_{2}^{*}\frac{d\psi_{2}}{dx}\bigg{)}+c.c\bigg{]}$
-- $\frac{1}{2}\bigg{|}\frac{d\psi_{1}}{dx}\bigg{|}^{2}-\frac{1}{2}\bigg{|}\frac{d\psi_{2}}{dx}\bigg{|}^{2}-\frac{\Delta}{2}|\psi_{1}|^{2}+\frac{\Delta}{2}|\psi_{2}|^{2}+\frac{\gamma_{0}}{2}|\psi_{1}|^{4}$
$+$ $\frac{{\gamma_{0}}}{2}|\psi_{2}|^{4}+{\beta_{0}}|\psi_{1}|^{2}|\psi_{2}|^{2}-V_{0}\cos(2k_{l}x)|\psi_{1}|^{2}$
-- $V_{0}\cos(2k_{l}x)|\psi_{2}|^{2}+\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{1}|^{4}$
$+$ $\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{2}|^{4}+\beta_{1}\cos(2k_{n}x)|\psi_{2}|^{2}|\psi_{1}|^{2}$
$+$ $\frac{i\alpha_{s}}{2}(\psi_{1}^{*}\frac{d\psi_{2}}{dx}+\psi_{2}^{*}\frac{d\psi_{1}}{dx})-\frac{i\alpha_{s}}{2}(\psi_{1}\frac{d\psi_{2}^{*}}{dx}+\psi_{2}\frac{d\psi_{1}^{*}}{dx}).$
A helicoidal spin-orbit coupled BEC supports freely moving bright solitons both in presence and absence of Zeeman splitting [32]. Thus, we adopt the following Gaussian-shaped trial solutions [17].
@@
$L=$displaystyle-- $\sum_{j\in\{1,2\}}N_{j}\bigg{[}\frac{d\phi_{j}}{dt}-k_{j}\frac{dx_{0}}{dt}+\frac{1}{4w^{2}}+\frac{k_{j}^{2}}{2}\bigg{]}$ (7)
$+$ $(k_{1}N_{1}-k_{2}N_{2})\beta$
$+$ $\frac{1}{2\sqrt{2\pi}w}(\gamma_{0}N_{1}^{2}+\gamma_{0}N_{2}^{2}+2\beta_{0}N_{1}N_{2})$
$+$ $\frac{e^{\frac{-k_{n}^{2}w^{2}}{2}}}{2\sqrt{2\pi}w}(\gamma_{1}N_{1}^{2}+\gamma_{1}N_{2}^{2}+2\beta_{1}N_{1}N_{2})\cos(2k_{n}x_{0})$
-- $V_{0}(N_{1}+N_{2})\,e^{-k_{l}^{2}w^{2}}\cos(2k_{l}x_{0})+\frac{\Delta}{2}(N_{2}-N_{1})$
-- $\alpha_{s}\sqrt{N_{1}N_{2}}(k_{1}+k_{2})e^{-k_{-}^{2}w^{2}}\cos(\phi_{1}-\phi_{2}).$
Making use of the Ritz optimization conditions: $\frac{\delta L}{\delta w}=0$, $\frac{\delta L}{\delta N_{1}}=0$, $\frac{\delta L}{\delta N_{2}}=0$, $\frac{\delta L}{\delta\phi_{1}}=0$, $\frac{\delta L}{\delta\phi_{2}}=0$, $\frac{\delta L}{\delta k_{1}}=0$ and $\frac{\delta L}{\delta k_{2}}=0$, we obtain the following coupled equations.
@@
with $\Lambda=N(\gamma_{0}-\beta_{0})/(\sqrt{2\pi})w$, $k_{\pm}=(k_{1}\pm k_{2})/2$, $G_{-}=\gamma_{1}-\beta_{1}$ and $G_{+}=\gamma_{1}+\beta_{1}$. Here, $\varphi=\phi_{1}-\phi_{2}$ and $Z=(N_{1}-N_{2})/N$ represent respectively phase difference and population imbalance between the two solitons. For a weak helicoidal gauge potential we take $k_{-}\approx\beta$. The coupled Eqs. (8)-(11) show that the rate of change of population imbalance depends directly on the initial phase difference and SOC parameters. Clearly, its dependence on the lattice and other parameters comes through $\varphi$. Therefore, it will be quite interesting to check how the population imbalance are affected by the different parameters of the system.
@@ III Helicoidal gauge potentials and JT oscillation
Let us consider that the coupled solitons are located at a particular position in the linear optical lattice (LOL). For these stationary solitons, the center of mass of the coupled system does not move in time and thus $\dot{x_{0}}=0$. In this case, the coupled Eqs. (8)-(11) can be reduced as follows.
@@
$G_{-}\geq\frac{4}{Z_{0}^{2}}\bigg{[}\sqrt{2\pi}w\frac{V_{0}k_{l}\sin(2k_{l}x_{0})}{\sin(2k_{n}x_{0})}e^{-\frac{w^{2}}{2}(2k_{l}^{2}+k_{n}^{2})}-\frac{G_{+}}{4}\bigg{]}$ (14)
with $10<Z_{0}<$ 1. Under this condition $k_{+}$ remains approximately constant with time. Therefore, Eqs. (12) and (13) can describe the variation of population imbalance with a change of relative phase between the two spin states in presence of both linear and nonlinear optical lattices (NOL).
@@
Figure 2: Left panel: Variation of population imbalance with $t$ in presence of nonlinear optical lattices for $\beta=1.6$. Here, solid blue and dashed black are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\varphi(0)=\pi/16$ respectively. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$ (solid blue) and $\beta=2$ (solid black) with $\varphi=\pi/8$. In both the panels, we take $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=1.05$, $V_{0}=2$, $k_{l}=5$, $k_{n}=1.8$, $\Delta=0$, $k_{0}=1$ and $\Lambda=0.04$.
Figure 3: Left panel:Variation of JT oscillation frequency with $s$. Here $G_{-}=1.05$. Right panel: Variation of $\delta\nu$ with $s$. In both the panels, we take $x_{0}=0.25$, $\Lambda=0.04$, $\alpha_{s}=4$ and $k_{n}=1.8$.
The effects of NOL on the JT oscillation in Halicoidal SOC-BEC is shown in Fig. 2. We see that the $Z(t)$ oscillates with time for different $\varphi(0)$ and the oscillations are quite similar to those observed in absence of NOLs. However, the frequency of oscillation ($\nu$) is larger than that found in the absence of NOL. This implies that the NOL causes acceleration in the dynamics of population imbalance.
@@
In writing Eq. (15) we have used $\ddot{Z}+\nu^{2}Z=0$. Understandably, $\nu$ in Eq. (15) does not contain $\Delta$ since its dependence on $\Delta$ is nonlinear. We plot $\nu$ for unequal strength of spin-orbit coupling ($\alpha$) and helicoidal gauge potential ($\beta$). We see that $\nu$ decreases with the increase of $s(\alpha-\beta)$. This indicates that the spin-orbit coupling and helicoidal gauge potential oppose each other in changing population imbalance (left panel). As a result, $\nu$ decreases with the increase of $s$. The frequency of JT oscillation in presence and absence of NOL are different (right panel of Fig. 3) and the difference decreases as $s$ increases.
@@
Variations of atomic densities with time in the two states of the system are shown in Fig. 4. It shows that the density of atomic population is changing between the two components (left and middle panel) periodically with time and the change of population imbalance is symmetric. This implies that maximum population in one component corresponds to minimum population of other component such that the total population remains constant with time (right panel).

### v4

@@ I Introduction
Josephson physics describes the tunneling phenomena that arise due to coupling between two macroscopically phase coherent quantum states such as superconductors separated by thin insulating layer. Due to this tunneling effect a constant supercurrent flows through the insulating layer in the case of d.c. Josephson effect while an oscillating supercurrent flows if a constant voltage is applied in the case of a.c. Josephson effect. Here the constant voltage induces temporal phase in suppercurrent. In weakly coupled Bose-Einstein condensates (BEC) separated by tunneling layer if a constant chemical potential difference is applied then the interplay between tunneling and interaction causes the oscillating atomic-supercurrent to flow and thus results an oscillation of atomic population difference between these two coupled BECs [1, 2]. This is termed as Josephson-type oscillation. If interaction dominates over the tunneling, a transition from Josephson oscillation to self-trapping occurs resulting the population imbalance to oscillate about a non-zero value. This tunneling oscillation can be explored in building quantum mechanical circuit to design qubit [3], SQUIDS [4] and to engineer quantum states [5].
A spin-orbit coupled (SOC) atomic BEC created by artificial gauge potentials offers a new possibility to study Josephson oscillation in pseudo-spin 1/2 systems. This type of oscillation in a weakly interacting BEC is a nonlinear generalization of the typical a.c. and d.c. Josephson oscillations. Thus it has received special attention and resulted several interesting experimental [6, 7, 8] and theoretical [9, 10, 11, 12, 13, 14, 15, 16, 17] investigations. Particularly, theoretical investigations in this direction include the studies on Josephson dynamics in two species BECs in double-well potential [10], internal Josephson oscillation between coupled solitons with time varying Raman frequency [17], effects of SOC on Josephson tunnelling between two annular condensates [18], Josephson oscillation of 2D chiral solitons [19] and Josephson-type oscillation in presence of optical lattices [20].
@@
Various techniques for generating artificial gauge potentials allow to tune strength and shape of spin-orbit coupling in BECs [26, 27, 28]. For the appropriate choice of gauge potential, ${\bf A}(x)=\boldsymbol{\sigma}\cdot\mathbf{n}(x)$ with $\mathbf{n}(x)=(\cos(2\beta x),\sin(2\beta x),0)$ and Pauli matrices ($\boldsymbol{\sigma}$)$=(\sigma_{1},\sigma_{2},\sigma_{3})$, one can induce $\pi/\beta$ periodic helicoidal non-uniform spin-orbit coupling in BECs. The periodic SO coupling has been considered to find its influence on the band structure, wave functions, and spin-density distribution of two-dimensional electron gas [29]. The so-called photonic graphene has been created through the generation of evanescently coupled helical waveguides where the propagation of light is described by vector potential [30].
@@
The helicoidal shaped vector potential ${\bf A}(x)$ introduces point translational symmetry in the system. However, one can switch to a rotating frame using an appropriate gauge transformation for a chosen gauge potential ${\bf A}(x)$ , where the point translational symmetry of the system transform into continuous translational symmetry. In this frame, the coefficients of the resulting equations become space-independent and the SOC becomes non-uniform. This non-uniform SOC is highly tunable and can increase the effects of intrinsic nonlinearity [26, 27, 28, 31, 32, 33]. A Bose-Einstein condensate with helicoidal SOC features many properties of integrable system such as stable propagation of bright solitons [34], Peregrine solitons [35], bright-dark Peregrine solitons [36], stripe solitons [37], multi-hump solitons [38] and kink solitons [39]. One of the important aspects of the soliton in helicoidal SOC-BECs is that the soliton with any peak amplitude is non-radiative and thus can propagate a large distance [34]. In a recent study it is shown that the helicoidal gauge potential can be used to manipulate modulational instability in BECs [40].
@@
Our objective in this paper is to study Josephson-type (JT) oscillation and self-trapping of atomic population between two solitons in helicoidal spin-orbit coupled quasi-one-dimensional (Q1D) BECs with optical lattices using an effective one-dimensional(1D) Gross-Pitaevskii equation(GPE). Particularly, we see that the atomic imbalance between the two solitons oscillates periodically and the frequency of oscillation depends sensitively on the strength of helicoidal gauge potential. The frequency decreases with the increase of helicoidal gauge potential. In presence of optical lattices, the frequency of oscillation further increases. We see that there exists a critical point for the observation of self-trapping. Interestingly, the critical point gets shifted in the parameter space with the change of helicoidal gauge potential and Zeeman-splitting. We note that there is no the so-called Rabi coupling, a linear coupling between the wave functions of the condensates, in helicoidal SOC-BECs. Here the two states are coupled through the gradient of the wave function i.e., helicoidal gauge potential. Therefore, we say that the oscillation of population imbalance can be affected by the helicoidal SOC. Recently, we consider uniform spin-orbit coupled BECs [41, 42, 43] in presence of nonlinear optical lattices (NOLs) and see that our approach can catch the effects of the NOLs and Rabi coupling on JT oscillation and self-trapping [20]. Here we extend the same to the case of non-uniform SOC-BECs with a view to extract the effect of helicoidal gauge potential on the system.
@@
In section II, we present a theoretical model to describe helicoidal SOC-BEC in optical lattices. Based on a variational approach, we derive equations for the center of mass, population imbalance and phase difference of the coupled condensates. In section III, we first fix the center of mass of soliton pair in a particular position in the linear optical lattice(LOL). Here the LOL acts as a trap. We then study the effects of periodic modulation of atomic interaction or the so-called nonlinear optical lattice (NOL) on the JT oscillations for different strengths of the helicoidal gauge potentials. In section IV, we find a condition for the observation of quantum mechanical self-trapping and its dependence on the helicoidal gauge potential and nonlinear optical lattice. Finally, we devote section V to summarize the results of our investigations.
@@ II General Formalism
$i\frac{\partial\psi_{j}}{\partial t}$ $=$ $-\frac{1}{2}\frac{\partial^{2}\psi_{j}}{\partial x^{2}}-i\beta(-1)^{j}\frac{\partial\psi_{j}}{\partial x}-i\alpha_{s}\frac{\partial\psi_{3-j}}{\partial x}+V_{ext}(x)\psi_{j}$ (1)
-- $\frac{\Delta}{2}(-1)^{j}\psi_{j}-(g_{1}|\psi_{j}|^{2}+g_{2}|\psi_{3-j}|^{2})\psi_{j},$
@@
$V_{ext}(x)=\frac{1}{2}{({\omega_{x}}/{\omega_{\perp}})^{2}}x^{2}+V_{0}\cos(2k_{l}x).$ (2)
Here, $\psi_{j}(x,t)$ $(j=1,2)$ gives the order parameter. The strengths of attractive intra- and inter-atomic interactions are represented by $g_{1}$ and $g_{2}$. Here we take same intra-atomic interaction for both the pseudo-spin states. In Eq. (1), the parameters $\alpha_{s}$, $\beta$ and $\Delta$ stand for the strengths of spin-orbit coupling, helicoidal gauge potential (HGP) and Zeeman splitting respectively. The external potential in Eq. (2) contains two terms, namely, harmonic trapping potential (first term) and linear optical lattice (second term) with wavenumber $k_{l}$ and strength $V_{0}$. We make Eq. (1) dimensionless by re-scaling the energy, length and time of the system in the units of $\hbar\omega_{\perp}$, $a_{\perp}=\sqrt{{\hbar}/{m\omega_{\perp}}}$ and $({\omega_{\perp}})^{-1}$ respectively. Here $\omega_{x}$ and $\omega_{\perp}$ stand for the angular frequencies of the harmonic trap in the axial and transverse directions.
With a view to consider a helicoidal spin-orbit coupled BEC trapped in linear optical lattices (LOL), we take $\omega_{\perp}\gg\omega_{x}$ such that the effects of harmonic trap is negligible. In addition, we introduce periodicity in the system by modulating the nonlinear mean-field interaction periodically, called nonlinear optical lattice (NOL). The NOLs are given by [44, 45, 46]
@@
where $k_{n}$ is the wavenumber of the NOLs with their strengths $\gamma_{1}$ and $\beta_{1}$ respectively. This lattice can be engineered by spatial periodic variation of either laser intensity in the vicinity of optical Feshbach resonance [47, 48] or magnetic field intensity in the vicinity of optically induced magnetic Feshbach resonance [49, 50]. Generally, the amplitude of periodic modulation is smaller than that of the unmodulated part of interaction such that $\gamma_{1}<\gamma_{0}$ and $\beta_{1}<\beta_{0}$. Understandably, $\gamma_{0}>\beta_{0}$ for miscible binary condensates. The NOLs are found very effective in controlling the dynamics and stability of matter-wave solitons in BECs [51, 52, 53, 54]. In order to proceed with the variational approach, we recast the Eqs. (1) and (2) in terms of the following Lagrangian density.
$\mathcal{L}$ $=$ $\sum_{j=1}^{2}\bigg{[}-\frac{1}{2}\bigg{|}\frac{d\psi_{j}}{dx}\bigg{|}^{2}+(-1)^{j}\frac{\Delta}{2}|\psi_{j}|^{2}+\frac{\gamma_{0}}{2}|\psi_{j}|^{4}$ (4)
$+$ $\bigg{(}\frac{i}{2}\bigg{(}\psi_{j}^{*}\frac{d\psi_{j}}{dt}+(-1)^{j}\beta\psi_{j}^{*}\frac{d\psi_{j}}{dx}+{\alpha_{s}}\psi_{j}^{*}\frac{d\psi_{3-j}}{dx}\bigg{)}+c.c\bigg{)}$
$+$ $\frac{\gamma_{1}}{2}\cos(2k_{n}x)|\psi_{j}|^{4}+\frac{\beta_{1}}{2}\cos(2k_{n}x)|\psi_{j}|^{2}|\psi_{(3-j)}|^{2}$
$+$ $\frac{\beta_{0}}{2}|\psi_{j}|^{2}|\psi_{(3-j)}|^{2}-V_{0}\cos(2k_{l}x)|\psi_{j}|^{2}\bigg{]}.$
A helicoidal spin-orbit coupled BEC supports freely moving bright solitons both in presence and absence of Zeeman splitting [34]. Thus, we adopt the following Gaussian-shaped trial solutions for bright solitons [17].
@@
Substituting Eq. (5) in Eq. (4) and then integrating with respect to $x$ from $\infty$- ∞ to +$\infty$+ ∞, we get the following Lagrangian.
$L$ $=$ $\!\sum_{j=1}^{2}\bigg{[}\!\!-N_{j}\bigg{(}\frac{d\phi_{j}}{dt}-k_{j}\frac{dx_{0}}{dt}+\frac{1}{4w^{2}}+\frac{k_{j}^{2}}{2}+(-1)^{j}\beta k_{j}\bigg{)}$ (6)
$+$ $\frac{1}{2\sqrt{2\pi}w}(\gamma_{0}N_{j}^{2}+\beta_{0}N_{j}N_{3-j})-\frac{V_{0}N_{j}}{e^{k_{l}^{2}w^{2}}}\,\cos(2k_{l}x_{0})$
$+$ $\frac{e^{\frac{-k_{n}^{2}w^{2}}{2}}}{2\sqrt{2\pi}w}(\gamma_{1}N_{j}^{2}+\beta_{1}N_{j}N_{3-j})\cos(2k_{n}x_{0})$
$+$ $(-1)^{j}\frac{\Delta N_{j}}{2}-\alpha_{s}\sqrt{N_{1}N_{2}}k_{j}e^{-k_{-}^{2}w^{2}}\cos(\varphi)\bigg{]}.$
Here $\varphi=\phi_{1}-\phi_{2}$. Making use of the Ritz optimization conditions: $\frac{\delta L}{\delta w}=0$, $\frac{\delta L}{\delta N_{1}}=0$, $\frac{\delta L}{\delta N_{2}}=0$, $\frac{\delta L}{\delta\phi_{1}}=0$, $\frac{\delta L}{\delta\phi_{2}}=0$, $\frac{\delta L}{\delta k_{1}}=0$ and $\frac{\delta L}{\delta k_{2}}=0$, we obtain the following coupled equations.
@@
with $\Lambda=N(\gamma_{0}-\beta_{0})/(\sqrt{2\pi})w$, $k_{\pm}=(k_{1}\pm k_{2})/2$, $G_{-}=\gamma_{1}-\beta_{1}$ and $G_{+}=\gamma_{1}+\beta_{1}$. Here, $\varphi$ and $Z=(N_{1}-N_{2})/N$ represent respectively phase difference and population imbalance between the two solitons. We know that $k_{-}$ is small since $k_{1}\approx k_{2}$ due to the assumption of same center of mass of both the solitons. Therefore, for weak helicoidal potential, we can take $k_{-}\approx\beta$. The coupled Eqs. (7)-(10) show that the rate of change of population imbalance depends directly on the initial phase difference and SOC parameters. Clearly, its dependence on the lattice and other parameters comes through $\varphi$. Therefore, it will be quite interesting to check how the population imbalance are affected by the different parameters of the system.
@@ III Helicoidal gauge potentials and JT oscillation
Let us consider that the coupled solitons are located at a particular position $(x_{0}={\rm const.})$ in the linear optical lattice (LOL). For these stationary solitons, the center of mass of the coupled system does not move in time and thus $\dot{x_{0}}=0$. In this case, the coupled Eqs. (7)-(10) can be reduced as follows.
@@
$G_{-}\geq\frac{4}{Z_{0}^{2}}\bigg{[}\frac{\sqrt{2\pi}wV_{0}}{e^{\frac{w^{2}}{2}(2k_{l}^{2}+k_{n}^{2})}}\frac{k_{l}\sin(2k_{l}x_{0})}{\sin(2k_{n}x_{0})}-\frac{G_{+}}{4}\bigg{]}.$ (13)
Here $Z_{0}$ stands for the initial population imbalance and it lies in the range $10<Z_{0}<$ 1. Under this condition $k_{+}$ remains approximately constant with time and thus we take $k_{+}=k_{0}$ in Eqs. (11) and (12). Therefore, Eqs. (11) and (12) can describe the variation of population imbalance with a change of relative phase between the two spin states in presence of both linear and nonlinear optical lattices (NOL).
@@
Figure 2: Left panel: Variation of population imbalance with t in presence of NOL for $\beta=1.6$. Here, solid-blue and dashed- black curves are drawn for the initial phase differences $\varphi(0)=\pi/8$ and $\varphi(0)=\pi/16$ respectively in presence of NOL while the dashed-blue and dotted-black curves are in absence of nonlinear optical lattices. Right panel: It gives $Z(t)$ versus $t$ for $\beta=1.6$ (solid-blue) and $\beta=2$ (solid-black) with $\varphi=\pi/8$. In both the panels, we take $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=1.05$, $V_{0}=2$, $k_{l}=5$, $k_{n}=1.8$, $\Delta=0$, $k_{0}=1$ and $\Lambda=0.04$.
Figure 3: Left panel: Variation of JT oscillation frequency with $s$. Here $G_{-}=1.05$. Right panel: Variation of difference ($\delta\nu$) between the frequencies in presence and absence of the NOL with $s$. In both the panels, we take $x_{0}=0.25$, $\Lambda=0.04$, $\alpha_{s}=4$, $k_{l}=5$ and $k_{n}=1.8$.
The effects of NOL on the JT oscillation in helicoidal SOC-BEC is shown in Fig. 2 taking $k_{n}<k_{l}$. Here the LOL is incommensurable with the NOL. We see that the $Z(t)$ oscillates with time for different $\varphi(0)$ and the oscillations are quite similar to those observed in absence of NOLs. However, the frequency of oscillation ($\nu$) is larger than that found in absence of NOL. This implies that the NOL causes acceleration in the dynamics of population imbalance. We have checked that the numerical value of $\nu$ decreases as $k_{n}\rightarrow k_{l}$.
@@
In writing Eq. (14) we have used $\ddot{Z}+\nu^{2}Z=0$. Understandably, $\nu$ in Eq. (14) does not contain $\Delta$ since its dependence on $\Delta$ is nonlinear. In order to understand the relative importance of spin-orbit coupling ($\alpha_{s}$) and helicoidal gauge potential ($\beta$) on JT oscillation, we define a parameter $s=\alpha_{s}-\beta$ and plot $\nu$ versus $s$ in Fig. 3 for non-uniform SOC by taking unequal values of $\alpha_{s}$ and $\beta$. We see that $\nu$ decreases with the increase of $s$. This indicates that the spin-orbit coupling and helicoidal gauge potential oppose each other in changing population imbalance (left panel). The frequency of JT oscillation with $s$ in presence and absence of NOL are different (right panel of Fig. 3) and their difference ($\delta\nu$) decreases as $s$ increases.
@@
Variations of atomic densities with time during JT oscillation in the two states of the system are shown in Fig. 4. It shows that the density of atomic population is changing between the two components (left and middle panel) periodically with time and the change of population imbalance is symmetric. This implies that maximum population in one component corresponds to minimum population of other component such that the total population remains constant with time (right panel).